C = 450 + 0.4Y
I = 350

G = 150
X = 70
Z = 35 + 0.1Y
T = 0.15Y
Yf = 1550



Q.2.6 Before the government decreased the tax rate, how much of government
spending was required to bring the economy to full employment?
1
Expert's answer
2020-12-03T07:11:33-0500

The equilibrium income level is:

"Y = C + I + G + NX = 450 + 0.4Y + 350 + 150 + 70 - (35 + 0.1Y) = 1,407.14."

The multiplier is:

"m = \frac {1} {1 - 0.4×(1 - 0.15) + 0.1} = 1.32."

The government spending should be increased by: "(1,550 - 1,407.14)\/1.32 = 108.23."